OpenStudy (fools101):

Imaginary Numbers Simplify i 9

OpenStudy (fools101):

Please help!

OpenStudy (pottersheep):

That is simplified isnt it?

OpenStudy (fools101):

oh also it i^9

OpenStudy (pottersheep):

ah. ok then so 1^2 = -1 right

OpenStudy (anonymous):

simplified version has to be i

OpenStudy (fools101):

Aww idk this kind of confusing me

OpenStudy (pottersheep):

(i^2)^4 . i (-1)^4 . i 1i

OpenStudy (pottersheep):

(i^2)^4 . i (-1)^4 . i 1i

OpenStudy (anonymous):

any variable equals 1 automatically

OpenStudy (pottersheep):

Do you get it?

OpenStudy (pottersheep):

Basically take out an i^2 first so that you work with -1. I.e i^4 = (i^2)^2 so now you have (-1)^2 = 1. Do you understand up to there?

OpenStudy (fools101):

KIND of so it 1? a little ...

OpenStudy (anonymous):

if there is no value to a variable it becomes a 1.

OpenStudy (fools101):

then the i Crossout ?

myininaya (myininaya):

\[i^9=i^{4 \cdot 2 +1}=i^{4 \cdot 2} \cdot i^1=(i^4)^2 \cdot i=(1)^2 \cdot i\] \[=1 \cdot i=i\]

OpenStudy (anonymous):

pattern is \[i^0=1\] \[i^1=i\] \[i^2=-1\] \[i^3=-i\] \[i^4=1\] and so on. so your real job is to take the integer remainder when you divide the power by 4. for example \[i^{103}=i^3=-i\]

OpenStudy (fools101):

ohhhh i see !!! kk thank u that make sense !!!

OpenStudy (anonymous):

yw